a speedy racer is drivint his car on a linear racing strip try to test out his cars specifications.he starts running the diagnostics while his car is travelling at a rate of 5 m/s. over the course of the next 5.4 seconds, he quadruples his speed and crosses the finish line. how far away was the finish line from where he started his diagnostic?
1
Expert's answer
2012-10-09T08:43:35-0400
The average speed of a car after running the diagnostics till crossing the finish line was

V = (4*5[m/s] + 5[m/s])/2 = 12.5[m/s].

So, the distance to the finish line from where he started his diagnostic was

D = 12.5[m/s]*5.4[s] = 67.5 metres.